# AutoGL: A Library for Automated Graph Learning

### 05 Jan 2024 v3

@@ 1 Introduction
Machine learning on graphs has drawn increasing popularity in the last decade [1]. For example, graph neural networks (GNNs) [2, 3] are the de facto standards and have shown successes in a wide range of applications such as recommendation [4, 5], traffic forecasting [6], physical simulations [7], geometric data analysis [8], bioinformatics [9], and combinatorial optimization [10].
@@ 2.9.5 Robust Graph Learning
  • Identity: keep the same adjacency matrix as in the previous layer, i.e., with adding robustness designs.
@@ 4 AutoGL-light
Since graph HPO aims to automatically optimize the hyper-parameters of graph machine learning models, as discussed in Section 2.6, we divide its designs into two parts: hyper-parameter (HP) space and search strategy. For the HP space, we support three types of search space: linear numerical search space, logarithmic numerical space, and categorical space. The former two are used to optimize numerical hyper-parameters, such as the number of layers, the dimensionality, the learning rate, etc., and the latter is used to optimize categorical hyper-parameters, such as the optimizer. In addition, users can also create their own hyper-parameter space by inheriting the base class. For the optimization algorithms, we support various black-box algorithms as in AutoGL (Section 2.6), such as grid search, random search, anneal algorithm, Bayes optimization, CMA-ES, MO-CMA-ES, Quasi-random, TPE, and AutoNE.
@@
To promote and showcase the flexibility and easy usage of AutoGL-light, we have implemented several state-of-the-art graph NAS methods, including out-of-distribution generalized graph NAS GRACES [80], large-scale graph NAS method GAUSS [81], and automated graph Transformer AutoGL [82]. Besides, we showcase AutoGL-light for handling various downstream graph tasks by applying AutoGL-light to bio-informatics using graph HPO and graph NAS, including ScGNN [83], MolCLR [84], and AutoGNNUQ [85]. These showcases were supported by GitLink Code Camp [^5] https://www.gitlink.org.cn/glcc.

### 05 Mar 2024 v4

@@ 1 Introduction
Machine learning on graphs has drawn increasing popularity in the last decade [1]. For example, graph neural networks (GNNs) [2, 3] are the de facto standards and have shown successes in a wide range of applications such as recommendation [4, 5], community detection [6], traffic forecasting [7], physical simulations [8], geometric data analysis [9], bioinformatics [10], and combinatorial optimization [11].
@@ 2.9.5 Robust Graph Learning
  • Identity: keep the same adjacency matrix as in the previous layer, i.e., without adding robustness designs.
@@ 4 AutoGL-light
Since graph HPO aims to automatically optimize the hyper-parameters of graph machine learning models, as discussed in Section 2.6, we divide its designs into two parts: hyper-parameter (HP) space and optimization algorithm. For the HP space, we support three types of search space: linear numerical search space, logarithmic numerical space, and categorical space. The former two are used to optimize numerical hyper-parameters, such as the number of layers, the dimensionality, the learning rate, etc., and the latter is used to optimize categorical hyper-parameters, such as the optimizer. In addition, users can also create their own hyper-parameter space by inheriting the base class. For the optimization algorithms, we support various black-box algorithms as in AutoGL (Section 2.6), such as grid search, random search, anneal algorithm, Bayes optimization, CMA-ES, MO-CMA-ES, Quasi-random, TPE, and AutoNE.
@@
To promote and showcase the flexibility and easy usage of AutoGL-light, we have implemented several state-of-the-art graph NAS methods, including out-of-distribution generalized graph NAS GRACES [81], large-scale graph NAS method GAUSS [82], and automated graph Transformer AutoGT [83]. Besides, we showcase AutoGL-light for handling various downstream graph tasks by applying AutoGL-light to bio-informatics using graph HPO and graph NAS, including ScGNN [84], MolCLR [85], and AutoGNNUQ [86]. These showcases were supported by GitLink Code Camp [^5] https://www.gitlink.org.cn/glcc. We aim to incorporate more advanced graph NAS such as [87, 88] and more applications such as [89, 90] in the future.